A grocery store bought milk for $2.80 per half gallon and stored it in two refrigerators. During the night, one refrigerator malfunctioned and ruined 13 half gallons. If the remaining milk is sold for $4.06 per half gallon, how many half gallons did the store buy if they made a profit of $63.14?
Answer by ikleyn(52781)

.
A grocery store bought milk for $2.80 per half gallon and stored it in two refrigerators.
During the night, one refrigerator malfunctioned and ruined 13 half gallons.
If the remaining milk is sold for $4.06 per half gallon,
how many half gallons did the store buy if they made a profit of $63.14?
~~~~~~~~~~~~~~~~~~~~~~

Let n be the number of half-gallons.


The bying cost was 2.80*n  dollars.

They sold (n-13) half-gallons, in all, and collected 4.06*(n-13) dollars.


The profit equation is

    4.06*(n-13) - 2.80n = 63.14.


Simplify and find n

    4.06n - 13*4.06 - 2.80n = 63.14

    4.06n - 2.80n = 63.14 + 13*4.06

         1.26n    =    115.92

             n    =    115.92/1.26 = 92.


ANSWER.  They bought 92 half-gallons.

Solved.